OpenStudy (anonymous):

What is the distance between (6, 2) and (-3, -2)? 9 3 5

OpenStudy (anonymous):

the other optoin is square root 97

OpenStudy (dmndlife24):

\[d = \sqrt{(x2-x1)^{2}+(y2-y1)^{2}}\]

pooja195 (pooja195):

the distance formula \[\huge~\rm~\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] We are trying to figure out the distance between the 2 point right? This formula will help us find the distance understood?

OpenStudy (anonymous):

i put it in and it said nonreal answer

pooja195 (pooja195):

\[\huge~\rm~\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] x1=? x2=? y1=? y2=?